OpenStudy (anonymous):

Solve. 3^(x – 5) = 9 a. x = –3 b. x = 2 c. x = 7 d. x = 8

OpenStudy (anonymous):

c

hero (hero):

@hossam1122 , don't just give an answer. Provide steps, explanations, and guidance.

OpenStudy (anonymous):

oh ok sorry about that so here 3^x-5=9 so 9 = 3^2 so 3^x-5=3^2 since they have same base 3 so x-5=2 x=7